OpenStudy (laronjaxon):

solve for y 5(y+4)-7y=8

OpenStudy (anonymous):

Distribute, then simplify. See what you come up with.

OpenStudy (laronjaxon):

I don't know what to distribute >.<

OpenStudy (goldphenoix):

Distribute 5 to y and 4.

OpenStudy (anonymous):

The term "5(y+4)" is really saying to you "5 multiplied by y AND 5 multiplied by 4." So start by distributing the 5 to the (y+4)

OpenStudy (goldphenoix):

In other word, 5 * y = ? 5 * 4 = ?

OpenStudy (laronjaxon):

5y+20

OpenStudy (goldphenoix):

Correct!

OpenStudy (goldphenoix):

Now you have 5y + 20 -7y = 8. Simplify it, like what jb said.

OpenStudy (laronjaxon):

would it be like 5y-7y which is 2?

OpenStudy (anonymous):

Remember, in "5y-7y" you aren't subtracting the "y", just the numbers. The y must stay.

OpenStudy (laronjaxon):

oh wait

OpenStudy (laronjaxon):

oh 2y?

OpenStudy (anonymous):

Close.

OpenStudy (laronjaxon):

y=2?

OpenStudy (goldphenoix):

That's 7y - 5y, not 5y - 7y.

OpenStudy (laronjaxon):

ugh >., soo that would make it 2y corrct :o? then what

OpenStudy (laronjaxon):

>.<*

OpenStudy (anonymous):

5-7 isn't 2. It's -2. Don't forget your signs! Always be careful.

OpenStudy (laronjaxon):

ahh :o! thank you both!

OpenStudy (anonymous):

Can you take it from there?

OpenStudy (laronjaxon):

I think so

OpenStudy (goldphenoix):

-2y = 8. You can do that right? If not, then ask jb. c:

OpenStudy (laronjaxon):

lol itll be -4 correct :p?

OpenStudy (anonymous):

Wait a sec, where did the 20 go? Haha. Don't forget the 20.

OpenStudy (goldphenoix):

Opps, my bad.

OpenStudy (anonymous):

Right now, you're at -2y + 20 = 8. Go from there.

OpenStudy (laronjaxon):

y=6

OpenStudy (anonymous):

Good job. That's correct.

OpenStudy (goldphenoix):

Good job jb for helping Laron! Proud of you.

OpenStudy (laronjaxon):

Thank you both ! I really appreciate it

OpenStudy (anonymous):

Anytime.